What is the initial pH of a 0.0025 M solution made pyrophosphoric acid. (This is an analytical chemistry question)
Answer #1

Ka of pyrophosphoric acid = 1.2 × 10−1

concentration = 0.0025 M

H4P2O7   ------------> H+ + H3P2O7-

0.0025                         0              0

0.0025 - x                     x              x

Ka = x^2 / 0.0025 - x

0.12 = x^2 / 0.0025 - x

x = 2.45 x 10^-3

pH = -log (2.45 x 10^-3)

pH = 2.61